Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%? Any help would be appreciated. Thank you very, very much.
Answer by checkley77(12844)

.09x+.08(9000-x)=770
.09x+720-.08x=770
.01x=770-720
.01x=50
x=50/.01
x=5000 is the amount invested @ 9%
9000-5000=4000 is the amount invested @ 8%
proof:
.09*5000+.08*4000=770
450+320=770
770=770